the perimeter of a rectanlge is 50 feet. the length is 5 more than the width. find the area of the rectangle

Answer by mananth(16949)

width =x
length = x+5
perimeter = 2*(L+W)
=2(x+x+5)
=2(2x+5)
=4x+10
4x+10=50
4x=50-10
4x=40
/4
x=10 feet the width
length = 10 +5 = 15 ft
Area = L*W
=10*15= 150 sq. ft